Canton Corp. produces a part using an expensive proprietary machine that can only be leased. The leasing company offers two cRequired A Required B Required C Required D What is the monthly break-even level assuming: Break-Even Level parts 1. 2. The uRequired A Required B Required C Required D At what volume would the operating profit be the same regardless of the lease optRequired A Required B Required C Required D Assume monthly volume of 29,500 units. What is the operating leverage assuming: (Required A Required B Required C Required D Assume monthly volume of 29,500 units. What is the margin of safety percentage as
Answer #1

solution (1) Unit rate lease of compute monthly break even level. Unit contribution on margin= unit selling price - unit vari607200 + 345 000 = 8280 (b) Let &- units under both produced the lease profit margin is options. Same 230% – 188 * - 607200 =Cu unit rate lease of prepare contribution margin income statement 6785000 407 1000 $2714000 Sales (29500X230) Less: variableoperating levarage = contribution margin operating income 3392500 2440300 = 1:39 ca unit rate leases Margin of safety = actuaflat rate lease r Margin of safety = actual sales-Break even sales = (29500x 230) – (828014230) - 6785000 - 1904400 = 4880600